OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that passes through (0, 5) and has a slope of −1? a--y = −x − 5 b--y = x + 5 c--y = −x + 5 d--y = x − 5

OpenStudy (yoloshroom):

so , slop intercept for is y=mx + b right? we know that M is so we can put that into the equation, y=-1x+b , and to find B we can substitute 0,5 into our equations, 5 being Y and 0 being X. So 5=-1(0)+b. Can you do the rest frm here?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 0}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies-1 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (yoloshroom):

damn that looks fancy. @jdoe0001

OpenStudy (jdoe0001):

hehe